# Sequential keypoint density estimator: an overlooked baseline of skeleton-based video anomaly detection

### v2

@@ 1 Introduction
In this manuscript, we propose SeeKer (Sequential Keypoint Density Estimator), a method for detecting anomalous human behaviour by predicting the conditional density of skeleton keypoints, as visualized in Figure 1. Given the preceding keypoints, SeeKer predict a multivariate normal distribution over possible locations for the subsequent keypoint. Applying this approach across the entire skeleton sequence results in an autoregressive factorization of the joint sequence density at the keypoint level. Consequently, SeeKer training corresponds to maximizing sequence likelihood. During inference, SeeKer identifies skeletons as anomalous if their constituent keypoints deviate from the predicted distributions. The corresponding skeleton anomaly score is a sum of keypoint likelihoods weighted by the keypoint extraction confidence from an off-the-shelf skeleton extractor. Thus, our final decision accounts for uncertainties across all pipeline components. Overall, SeeKer presents the following key novelties to skeleton-based anomaly detection: autoregressive density factorization at the keypoint level, interpretable compositional anomaly scoring, and principled inclusion of keypoint detection uncertainty. Experimental evaluation shows that SeeKer achieves six percentage points absolute improvement over the best previous method on the UBnormal dataset and five percentage points absolute improvement on the MSAD-HR dataset in terms of AUROC.
@@ 2 Related work
Video anomaly detection. Video anomaly detection [43, 30] focuses on identifying anomalous frames within video sequences. This task is particularly relevant in surveillance [48] where anomalies may indicate accidents or emergencies [34]. Early approaches to frame-based video anomaly detection map input frames into a shared feature space [14, 2] and apply decision rules that differentiate between normal and abnormal patterns [25]. More recent approaches [16, 19, 8] focus on bounding box detections and recognize anomalies using self-supervised features [20, 53]. Other methods leverage failed optical flow predictions [36], errors in denoising of diffusion model [55], and predictions from cooperatively trained deep models [59]. Furthermore, recent methods use exciting density estimation techniques such as denoising score matching [42]. However, these RGB-based methods may raise ethical concerns due to potential appearance-based biases and difficult anonymization. SeeKer builds atop skeleton sequences, and thus avoids such concerns.
@@
Autoregressive data likelihood factorization. Autoregressive factorization of data likelihood emerged as a foundational approach across numerous machine learning applications with dominant temporal domain [5, 12]. Such factorization is conveniently modeled using sliding window based approaches [52, 6], recurrent neural networks [46, 24] or transformer decoders [51, 9] with causal self-attention layers [50, 32]. SeeKer relies on autoregressive factorization of skeleton sequence density.
@@ 3 Skeleton anomaly detection by predicting probable locations of subsequent keypoint
The core approach behind SeeKer is to predict a conditional distribution of probable keypoint locations based on the preceding keypoints in the skeleton sequence. Leveraging this design in conjunction with the compositional nature of skeletons provides a highly interpretable skeleton anomaly score that aggregates per-keypoint likelihoods.
@@ 3.1 Autoregressive density of skeleton sequences
We derive the conditional per-keypoint density from the joint density of a skeleton sequence with a top-down approach. The joint density of a skeleton sequence $\mathbf{X}$ can be conveniently factorized by autoregressive modelling [21, 7] in order to capture skeleton motion:
@@
Here, the conditional likelihood of a skeleton $X_{t}$ is determined from the skeletons from the preceding $t-1$ time steps, denoted by $\mathbf{X}_{<t}$. Given skeleton sequences of arbitrary length, we adopt a siding-window approach with a fixed-size time window. In practice, this means that we can faithfully approximate $p_{\theta}(X_{t}|\mathbf{X}_{<t})$ using only past $\Delta$ skeletons, i.e., $p_{\theta}(X_{t}|\mathbf{X}_{\Delta})$, where $\Delta$ is sufficiently large to capture crucial temporal dependencies.
@@
We condition the density of $n$-th keypoint at timestep $t$ on the previous $n-1$ keypoints of the current timestep, as well as $N\cdot\Delta$ keypoints from the preceding timesteps. This formulation requires a fixed ordering of keypoints across skeletons. However, as experimentally confirmed, keypoint ordering does not affect the model’s expressiveness [40].
We further define per-keypoint conditional likelihoods as multivariate normal distributions with mean $\boldsymbol{\mu}_{\theta}$ and covariance $\Sigma_{\theta}$ computed from the preceding keypoints:
@@
Here, both $\boldsymbol{\mu}_{\theta}$ and $\Sigma_{\theta}$ are predicted from the preceding keypoints in the current timestep $X_{t,<n}$ and the context $\mathbf{X}_{\Delta}$ using a deep autoregressive model with tuneable parameters $\theta$. Altogether, we can now express the log-likelihood of the entire sequence (1) in terms of the keypoint likelihoods:
@@
[…]
@@
Figure 3: Graphical representation of the causality constraints in our autoregressive factorization. The first (light-blue) output depends only on the context, while the yellow output depends on all keypoints from the context and its predecessors (light-blue and red keypoints) from the target.
The described causal constraint can be enforced using different architectures [24, 21, 51]. Our experimental evaluation revealed that the relatively simple masked fully connected model [21] performs best for our task, which can be attributed to the uniformity of the skeleton sequences. Our model differs from [21] by following the natural ordering of inputs while allowing neighboring inputs (e.g., keypoint coordinates) to be codependent rather than strictly causal. Thus, we enforce causality through fixed blockwise triangular masks. We revisit this architecture in Appendix B. Next, we elaborate how SeeKer accurate anomaly detection.
@@ 3.2 Composite anomaly score
SeeKer expresses the density of a skeleton sequence as aggregate key-point density (Eq. 2). Thus, observing a keypoint at a probable location indicates normal (expected) behavior. Contrary, observing the keypoint at an unprobable location signals abnormal behavior. We define the anomaly score of the skeleton $X_{t}$ given the context $\mathbf{X}_{\Delta}$ as a joint density of its keypoints:
@@
The anomaly score (7) assumes that the entire skeleton $X_{t}$ is accurately extracted from the corresponding scene $\mathcal{I}_{t}$. However, occlusions and faulty skeleton detectors may produce errors. Fortunately, most detectors provide a confidence score for each keypoint [15]. Thus, we introduce the detection confidence into our anomaly score as:
@@ 4 Experimental setup
Datasets. We consider three publicly available datasets for unsupervised video anomaly detection in skeleton sequences. UBnormal [3] is a synthetic dataset with 268 training, 64 validation, and 211 test video sequences. The inlier sequences contain actions such as walking while texting and talking to others. Anomalous actions include unexpected behaviour, such as stealing or fighting. This is the most relevant dataset for our task due to highly accurate labels. ShanghaiTech [34] consists of 330 training and 107 test videos of real-world scenes. Walking, standing, or sitting are considered as inlier actions, while cycling, skateboarding, and fighting are anomalous. Furthermore, both datasets contain anomalous events that are not human-related (e.g. car accident or fire). Consequently, we also consider dataset subsets that contain only human-related anomalies as suggested in [16, 45] refered to as UBnormal-HR and ShanghaiTech-HR. The MSAD [62] dataset includes videos captured across diverse locations and various viewpoints, ensuring a wide range of perspectives and conditions, making it the most challenging benchmark to date. MSAD contains 720 videos in total, some extending up to 6K frames, and features 55 distinct abnormal events. We focus on the human-related subset with 20 distinct abnormal actions, MSAD-HR, as the full dataset track includes scenarios and events related to industrial automation and manufacturing without any human involvement. Note that non-human related anomalies on UBnormal and ShanghaiTech still involve humans, making them detectable through human reactions to events like fire or unusual vehicle behavior. Previous work [23] noticed that contemporary skeleton extractors [15] and trackers [54] fail to produce satisfactory outputs on some datasets (UCFCrime [48] and Avenue [37]). Thus, we skip these datasets in our experimental evaluation.
@@
We predict the per-keypoint distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$ with deep autoregressive models that we configure for skeleton sequences. We consider causally masked fully connected models [21], transformers [51], and recurrent models. We train the causal fully connected model for 20 epochs on UBnormal and MSAD-HR and 10 epochs on ShanghaiTech with the Adam optimizer, learning rate $10^{-3}$ and batch size 256. Since UBnormal is the only dataset with a validation split, we validate all hyperparameters on the UBnormal validation subset and reuse them for ShanghaiTech and MSAD-HR. During inference, SeeKer considers $T$ preceding frames in a sliding-window manner. At the beginning of a sequence, we backtrack the score of the $T$-th frame to the first $T-1$ frames. Following [23, 16, 42], we finally apply Gaussian smoothing of the computed anomaly scores. However, we apply smoothing for each person separately before aggregating the per-frame anomaly score with $\sigma$ = 10. All reported results are averaged over three runs, and variance is consistently below $0.05$ pp. Our code is publicly available here [^1] https://github.com/adelic99/seeker.
@@ 5.2 Qualitative performance
Figure 4 compares SeeKer and the best skeleton-oriented baseline STG-NF [23] on an anomalous sequence from the ShanghaiTech dataset (bike riding event in the pedestrian area). The x-axis represents time, while the y-axis indicates anomaly scores. We plot both, the raw anomaly scores and the anomaly scores smoothed with a 1D Gaussian filter. Red-shaded regions mark frames labeled as anomalous. SeeKer accurately identifies the anomalous frames, while STG-NF fails to detect anomalies at the temporal borders of the abnormal action. However, as shown later in Table 7, smoothing contributes to the final AUROC accomplished by STG-NF [23] for even 4pp since it cancels out errors at the temporal borders. Previous work [42] observes that the ShanghaiTech dataset contains noisy labels in the context of skeleton-based anomaly detection. For example, a bicycle wheel may enter the camera view before the rider, which results in a later than anticipated alarm of skeleton-based approaches. Nonetheless, SeeKer attains competitive results on the ShanghaiTech dataset, despite these disadvantages.
@@
Figure 6: Interpretability of SeeKer demonstrated on an anomalous action from the UBnormal dataset (running). SeeKer anomaly score aggregates per-keypoint likelihoods denoted with varying color intensities. Each keypoint contribution to the final score changes over time.
The SeeKer anomaly score (8) is highly interpretable, as it consolidates keypoint likelihoods into skeleton-level anomaly scores while operating in the original 2D space. Figure 6 illustrates the interpretability by tracking the log-likelihood of each keypoint over time on the UBnormal dataset. At each timestep, different keypoints contribute to the skeleton’s abnormality, highlighted by varying color intensities. Notably, this level of analysis is not feasible with previous skeleton-based methods.
@@ 5.3 Discussion
For all experiments in the discussion, we report AUROC on the UBnormal and ShanghaiTech test.
@@ 7 Acknowledgements
This work has been supported by Croatian Recovery and Resilience Fund - NextGenerationEU (grant C1.4 R5-I2.01.0001).
@@ B.1 Causal fully connected model
[…]
@@ B.2 Causal transformer
[…]
@@ Appendix C Additional discussion
C.3 On computational requirements.

### v3

@@ 1 Introduction
In this manuscript, we propose SeeKer (Sequential Keypoint Density Estimator), a method for detecting anomalous human behaviour by predicting the conditional density of skeleton keypoints, as visualized in Figure 1. Given the preceding keypoints, SeeKer predicts a multivariate normal distribution over possible locations for the subsequent keypoint. Applying this approach across the entire skeleton sequence results in an autoregressive factorization of the joint sequence density at the keypoint level. Consequently, SeeKer training corresponds to maximizing sequence likelihood. During inference, SeeKer identifies skeletons as anomalous if their constituent keypoints deviate from the predicted distributions. The corresponding skeleton anomaly score is a sum of keypoint likelihoods weighted by the keypoint extraction confidence from an off-the-shelf skeleton extractor. Thus, our final decision accounts for uncertainties across all pipeline components. Overall, SeeKer presents the following key novelties to skeleton-based anomaly detection: autoregressive density factorization at the keypoint level, interpretable compositional anomaly scoring, and principled inclusion of keypoint detection uncertainty. Experimental evaluation shows that SeeKer achieves six percentage points absolute improvement over the best previous method on the UBnormal dataset and five percentage points absolute improvement on the MSAD-HR dataset in terms of AUROC.
@@ 2 Related work
Video anomaly detection. Video anomaly detection [43, 30] focuses on identifying anomalous frames within video sequences. This task is particularly relevant in surveillance [48] where anomalies may indicate accidents or emergencies [34]. Early approaches to frame-based video anomaly detection map input frames into a shared feature space [14, 2] and apply decision rules that differentiate between normal and abnormal patterns [25]. More recent approaches [16, 19, 8] focus on bounding box detections and recognize anomalies using self-supervised features [20, 53]. Other methods leverage failed optical flow predictions [36], errors in denoising of diffusion models [55], and predictions from cooperatively trained deep models [59]. In addition, there are exciting density estimation techniques such as denoising score matching [42]. However, these RGB-based methods may raise ethical concerns due to potential appearance-based biases and difficult anonymization. SeeKer builds atop skeleton sequences, and thus avoids such concerns.
@@
Autoregressive data likelihood factorization. Autoregressive factorization of data likelihood emerged as a foundational approach across numerous machine learning applications with a dominant temporal domain [5, 12]. Such factorization is conveniently modeled using sliding window based approaches [52, 6], recurrent neural networks [46, 24] or transformer decoders [51, 9] with causal self-attention layers [50, 32]. SeeKer relies on autoregressive factorization of skeleton sequence density.
@@ 3 Skeleton anomaly detection by predicting probable locations of subsequent keypoint
The core approach behind SeeKer is to predict a conditional distribution of probable keypoint locations based on the preceding keypoints in the skeleton sequence, as visualized in Figure 2. Leveraging this design in conjunction with the compositional nature of skeletons provides a highly interpretable skeleton anomaly score that aggregates likelihoods estimated at the keypoint level.
@@ 3.1 Autoregressive density of skeleton sequences
We derive the conditional per-keypoint density from the joint density of a skeleton sequence with a top-down approach. The joint density of a skeleton sequence $\mathbf{X}$ can be conveniently factorized by autoregressive modeling [21, 7] in order to capture skeleton motion:
@@
Here, the conditional likelihood of a skeleton $X_{t}$ is determined from the skeletons from the past t1t-1 time steps, denoted by $\mathbf{X}_{<t}$. Given skeleton sequences of arbitrary length, we adopt a siding-window approach with a fixed-size time window. In practice, this means that we can faithfully approximate $p_{\theta}(X_{t}|\mathbf{X}_{<t})$ using only past $\Delta$ skeletons, i.e., $p_{\theta}(X_{t}|\mathbf{X}_{\Delta})$, where $\Delta$ is sufficiently large to capture crucial temporal dependencies.
@@
We condition the density of $n$-th keypoint at timestep $t$ on the preceding n1n-1 keypoints of the current timestep, as well as $N\cdot\Delta$ keypoints from the past timesteps, as ilustrated on Figure 3. This formulation requires a fixed ordering of keypoints across skeletons. However, as experimentally confirmed later, keypoint ordering does not affect the model’s expressiveness [40].
We further define per-keypoint conditional likelihoods as multivariate normal distributions with mean $\boldsymbol{\mu}_{\theta}$ and covariance $\Sigma_{\theta}$ computed from the previous keypoints:
@@
Both $\boldsymbol{\mu}_{\theta}$ and $\Sigma_{\theta}$ are predicted from the preceding keypoints in the current timestep $X_{t,<n}$ and the past keypoints $\mathbf{X}_{\Delta}$ using a deep autoregressive model with tuneable parameters $\theta$. Altogether, we can now express the log-likelihood of the entire sequence (1) in terms of the keypoint likelihoods:
@@
Figure 3: We autoregressively predict distributions of possible keypoint locations for the current skeleton $X_{t}$. For instance, the distribution of the current keypoint $X_{t,n}$ (denoted in red) is estimated from all past keypoints $X_{\Delta}$ and the preceding keypoints $X_{t,<n}$, all denoted in blue. Dark keypoints are masked due to autoregressive likelihood factorization.
@@
$\theta_{\text{MLE}}=\underset{\theta}{\text{argmax}}\,L(\theta;\mathcal{D}),L(\theta;\mathcal{D})=-\sum_{\mathbf{X}\in\mathcal{D}}\ln p_{\theta}(\mathbf{X}).$ (5)
@@
Figure 4: Graphical representation of the causality constraints in autoregressive factorization. The first (blue) output depends only on past keypoints, while the green output depends on all past keypoints and preceding blue and red keypoints from the target.
The described causal constraint can be enforced using different architectures [24, 21, 51]. Our experimental evaluation revealed that the relatively simple masked fully connected model [21] performs best for our task, which can be attributed to the uniformity of the skeleton sequences. Our model differs from [21] by following the natural ordering of inputs while allowing neighboring inputs (e.g., keypoint coordinates) to be codependent rather than strictly causal. Thus, we enforce causality through fixed blockwise triangular masks. We revisit this architecture in Appendix B.
@@ 3.2 Composite anomaly score
SeeKer expresses the density of a skeleton sequence as aggregate key-point density (Eq. 2). Thus, observing a keypoint at a probable location indicates normal (expected) behavior. Contrary, observing the keypoint at an improbable location signals abnormal behavior. We define the anomaly score of the skeleton $X_{t}$ given the context $\mathbf{X}_{\Delta}$ as a joint density of its keypoints:
@@
The anomaly score (7) assumes that the entire skeleton $X_{t}$ is accurately extracted from the corresponding scene $\mathcal{I}_{t}$. However, occlusions and faulty skeleton detectors may produce errors. Fortunately, most detectors provide a confidence score for each keypoint [15]. Consequently, we introduce the detection confidence into our anomaly score as:
@@ 4 Experimental setup
Datasets. We consider three publicly available datasets for unsupervised video anomaly detection in skeleton sequences. UBnormal [3] is a synthetic dataset with 268 training, 64 validation, and 211 test video sequences. The inlier sequences contain actions such as walking while texting and talking to others. Anomalous actions include unexpected behaviour, such as stealing or fighting. This is the most relevant dataset for our task due to highly accurate labels. ShanghaiTech [34] consists of 330 training and 107 test videos of real-world scenes. Walking, standing, or sitting are considered as inlier actions, while cycling, skateboarding, and fighting are anomalous. Furthermore, both datasets contain anomalous events that are not human-related (e.g. car accident or fire). Consequently, we also consider dataset subsets that contain only human-related anomalies as suggested in [16, 45] refered to as UBnormal-HR and ShanghaiTech-HR. The MSAD [62] dataset includes videos captured across diverse locations and various viewpoints, ensuring a wide range of perspectives and conditions, making it the most challenging benchmark to date. MSAD contains 720 videos in total, some extending up to 6K frames, and features 55 distinct abnormal events. We focus on the human-related subset with 20 distinct abnormal actions, MSAD-HR, as the full dataset track includes scenarios and events related to industrial automation and manufacturing without any human involvement. Note that non-human related anomalies on UBnormal and ShanghaiTech still involve humans, making them detectable through human reactions to events like fire or unusual vehicle behavior. Previous work [23] noticed that contemporary skeleton extractors [15] and trackers [54] fail to produce satisfactory outputs on some datasets (UCFCrime [48] and Avenue [37]). Hence, we skip these datasets in our experiments.
@@
We predict the per-keypoint distribution parameters $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$ with deep autoregressive models that we configure for skeleton sequences. We consider causally masked fully connected models [21], transformers [51], and recurrent models. We train the causal fully connected model for 20 epochs on UBnormal and MSAD-HR and 10 epochs on ShanghaiTech with the Adam optimizer, learning rate $10^{-3}$ and batch size 256. Since UBnormal is the only dataset with a validation split, we validate all hyperparameters on the UBnormal validation subset and reuse them for ShanghaiTech and MSAD-HR. During inference, SeeKer considers $T$ preceding frames in a sliding-window manner. At the beginning of a sequence, we backtrack the score of the $T$-th frame to the first T1T-1 frames. Following [23, 16, 42], we finally apply Gaussian smoothing of the computed anomaly scores. However, we apply smoothing for each person separately before aggregating the per-frame anomaly score with $\sigma$ = 10. All reported results are averaged over three runs, and variance is consistently below $0.05$ pp. Our code is publicly available [^1] https://github.com/adelic99/seeker.
@@ 5.2 Qualitative performance
The SeeKer anomaly score (8) is highly interpretable, as it consolidates keypoint likelihoods into skeleton-level anomaly scores while operating in the original 2D space. Figure 5 shows the difference in predictions during normal and abnormal behavior. The observations align closely with predictions of possible keypoint locations at normal timesteps. At abnormal timesteps, the predictions tend to follow normal motion patterns and deviate significantly from actual observations, resulting in high anomaly scores. In addition, Figure 6 emphasizes the interpretability of per-keypoint contributions. At each timestep, keypoints contribute with different intensity to the abnormality of the skeleton, as visualized through color variation. Notably, this level of analysis is not feasible with previous methods.
Figure 5: Qualitative performance of SeeKer on an example from UBnormal. During normal behavior, the observed keypoints emerge at anticipated locations. Abnormal behavior causes significant deviation from the expected keypoint locations, which enables seamless anomaly detection.
Figure 6: Interpretability of SeeKer shown on an anomalous action from UBnormal (running). SeeKer anomaly score aggregates per-keypoint likelihoods denoted with varying color intensities. Each keypoint contribution to the final score changes over time.
Figure 7 compares SeeKer and the best skeleton-oriented baseline STG-NF [23] on an anomalous sequence from ShanghaiTech (bike riding in the pedestrian area). We plot both the raw anomaly scores and the anomaly scores smoothed with a 1D Gaussian filter. The x- and y-axes represent time and anomaly score, respectively, and red-shaded regions mark frames labeled as anomalous. SeeKer accurately identifies the anomalous frames, while STG-NF fails to detect anomalies at the temporal borders of the abnormal action. Prior work [42] notes that ShanghaiTech contains noisy labels for skeleton-based anomaly detection. For instance, a bicycle wheel may appear before the rider, delaying detection by skeleton-based methods.
@@ 5.3 Discussion
For all experiments in the discussion, we report AUROC on the UBnormal and ShanghaiTech test subsetss.
@@ 7 Acknowledgements
This work has been supported by Croatian Recovery and Resilience Fund - NextGenerationEU (grant C1.4 R5-I2.01.0001). The authors declare no competing interests.
@@ B.1 Causal fully connected model
$\text{C-FC${}_{l}$}(\mathbf{x})=(\mathbf{W}\odot\mathbf{M})\cdot\mathbf{x}+\mathbf{b}.$ (16)
@@ B.2 Causal transformer
$\text{C-SA}(\mathbf{X})=\left(\sigma(\mathbf{X}\mathbf{W}_{q}\mathbf{W}_{k}^{\top}\mathbf{X}^{\top}/\sqrt{d})\odot\mathbf{M}^{\text{SA}}\right)\mathbf{X}\mathbf{W}_{v}.$ (22)
@@ Appendix C Additional discussion
C.3 On non-human related anomalies.
C.4 On computational requirements.